simple mixtures

addition of 100g of a compound to 750g of cci4 lowered the freezing point of the solvent by 10.5 k.calculate the molar mass of the compound
Answer #1
Step1 Let M be molar mass ; Moles of solute =100/M

Step2 Moles of solute per kg solvent = (100/M)x(1000/750)= 400/3M ; Kf =30.0 for CCl4

Step3 Kf * m = 10.5 ; 30x400/3M= 10.5 ; M=380.95
answered by: sarena